### v1

# Learning Constraints from Stochastic Partially-Observed Closed-Loop Demonstrations

Abstract
We present an algorithm for learning unknown parametric constraints from locally-optimal input-output trajectory data. We assume that the given data is generated by demonstrators with stochastic nonlinear dynamics who execute a state or output feedback law to robustly satisfy the constraints despite worst-case dynamics and output noise. We encode the Karush-Kuhn-Tucker (KKT) conditions of this robust optimal output feedback control problem within a feasibility problem to recover constraints consistent with the local optimality of the demonstrations. We prove that our constraint learning method (i) accurately recovers the demonstrator’s state or output feedback policy, and (ii) conservatively estimates the set of all state or output feedback policies that ensure constraint satisfaction despite worst-case noise realizations. Moreover, we perform sensitivity analysis, proving that when demonstrations are corrupted by transmission error, the inaccuracy in the learned state or output feedback law scales linearly in the error magnitude. Our method accurately recovers unknown constraints from simulated noisy, closed-loop demonstrations generated using dynamics, both linear and nonlinear, (e.g., unicycle and quadrotor) and a range of state and output feedback mechanisms.
I Introduction
In safety-critical robotics, autonomous agents must infer and enforce hard constraints to guarantee safe operation, such as for collision avoidance in crowded environments. Learning from demonstration (LfD) enables robots to recover unknown parametric constraints from trajectory data [1, 2, 3]. The key idea is that locally-optimal demonstrations generated from a constrained optimal control problem satisfy an associated set of Karush-Kuhn-Tucker (KKT) conditions, which can be inverted to recover constraints via inverse optimal control (IOC). Prior work has shown this is possible for open-loop state and control demonstrations under deterministic dynamics [1, 4]. In practice, however, agents operate with state uncertainty and stochastic dynamics. Thus, instead of open-loop plans, a demonstrator must apply an output feedback control policy to partial, noisy observations. Existing methods cannot recover constraints from such input-output data.
To address this gap, we make the following contributions. We present a novel method to infer a priori unknown parametric constraints from input-output data generated under linear time-varying (LTV) stochastic dynamics with partial observations and output noise. Our method extends the IOC-based constraint learning method in [1] to the partial and noisy observation setting (Sec. III). We evaluate our method on linear and nonlinear (e.g., unicycle, quadrotor) dynamics and a variety of feedback controllers, successfully recovering constraints from the generated demonstrations (Sec. IV).
Our work draws from and contributes to IOC-based LfD for cost and constraint inference. Specifically, [5, 6, 2] introduce IOC methods assuming known constraints, [7, 8] enable cost inference via IOC, while [1, 9, 10] present IOC methods for constraint learning. Unlike [9], which addresses stochastic demonstrations in discrete spaces, our method recovers constraints with provable accuracy guarantees in continuous domains. Compared to [10], which assumes demonstrators with partial obstacle knowledge but full state information, we consider partial state observability. Further, while [1] use deterministic trajectories, we learn constraints from demonstrations generated via output feedback with noisy, partial observations. On a technical level, we leverage system-level synthesis (SLS) [11, 12, 13] to encode output feedback policies for input-output demonstrations. Whereas prior SLS work focuses on controller design, we apply SLS to constraint inference from noisy, partially observed data.
Notation: Given $x_{0},\cdots,x_{T}\in\mathbb{R}^{n}$ and indices $t_{1},t_{2}\in\{0,\ldots,T\}:=[0,T]$ with $t_{1}\leq t_{2}$, let $x_{t_{1}:t_{2}}:=(x_{t_{1}},\cdots,x_{t_{2}})\in\mathbb{R}^{n(t_{2}-t_{1}+1)}$. Given $M_{1},\cdots,M_{K}\in\mathbb{R}^{m\times n}$, we denote by $M:=\text{diag}\{M_{1},\cdots,M_{K}\}\in\mathbb{R}^{mK\times nK}$ the block diagonal matrix with $M_{k}$ as the $k$-th block, $\forall k\in\{1,\ldots,K\}:=[K]$. For any $m,n\in\mathbb{N}$, we denote the $m\times m$ identity matrix by $I_{m}$ and the $m\times n$ zero matrix by $O_{m\times n}$, with subscripts omitted when dimensions are clear from context. Given $A\in\mathbb{R}^{m\times n}$, we denote by $A^{\dagger}\in\mathbb{R}^{n\times m}$ the Moore-Penrose pseudoinverse of $A$.
II Problem Formulation
We formulate the robust optimal control problem used to generate demonstration data (Sec. II-A), present system level synthesis (SLS) for output feedback (Sec. II-B), discuss data generation (Sec. II-C), and state our problem (Sec. II-D).
II-A Robust Optimal Control Problem Formulation
Consider a stochastic dynamical system $f_{t}:\mathbb{R}^{n}\times\mathbb{R}^{n_{i}}\rightarrow\mathbb{R}^{n}$ with additive dynamics and output noise, and initial condition fixed at $x_{0}=\bar{x}$ as shown below: [^1] Uncertainty in $x_{0}$ can be modeled by incorporating it into the definition of the dynamics noise; we do so in our experiments. For details, see [12].
$x_{t+1}$ $=f_{t}(x_{t},u_{t})+w_{t},\hskip 14.22636pt$ $\forall t\in[0,T-1],$ (1a)
$y_{t}$ $=C_{t}x_{t}+e_{t},$ $\forall t\in[0,T].$ (1b)
Above, $x_{t}\in\mathbb{R}^{n}$, $u_{t}\in\mathbb{R}^{n_{i}}$, $w_{t}\in\mathbb{R}^{n}$, $y_{t}\in\mathbb{R}^{n_{o}}$, and $e_{t}\in\mathbb{R}^{n_{o}}$ denote the state, input, dynamics noise, output, and output noise at time $t\in[0,T]$, respectively. For convenience, let $\textbf{x}:=x_{0:T}\in\mathbb{R}^{n(T+1)}$, $\textbf{u}:=u_{0:T-1}\in\mathbb{R}^{n_{i}T}$, $\textbf{w}:=w_{0:T-1}\in\mathbb{R}^{nT}$, and $\textbf{y}:=y_{0:T},\textbf{e}:=e_{0:T-1}\in\mathbb{R}^{n_{o}T}$. We assume there exist known, compact sets $W\subset\mathbb{R}^{n}$ and $E\subset\mathbb{R}^{n_{o}}$ such that $w_{t}\in W$ and $e_{t}\in E$ for each $t\in[T]$. Note that $W$ and $E$ can be estimated from data, e.g., via [14].
To ease exposition, unless noted otherwise, we assume linear time-varying (LTV) dynamics, i.e., $f_{t}(x_{t},u_{t})=A_{t}x_{t}+B_{t}u_{t}$. We add clarifying remarks when a result holds for LTV but not for general nonlinear dynamics (Remarks 1, 2, 6).
Let $C_{k}$, $C_{\urcorner{k}}(\theta^{\star})\subset\mathbb{R}^{n(T+1)+n_{i}T}$ respectively denote a known and an unknown, parameterized constraint set imposed on the state and control trajectory $(\textbf{x},\textbf{u})$, where $\theta^{\star}\in\mathbb{R}^{d}$ denotes a ground truth parameter vector a priori unknown to the constraint learner. Concretely, $C_{k}$ and $C_{\urcorner{k}}(\theta^{\star})$ are respectively defined by inequality constraints $g_{\beta,k}:\mathbb{R}^{n(T+1)+n_{i}T}\rightarrow\mathbb{R}^{n_{k}},\forall\beta\in[n_{k}]$, and $g_{\beta,{\urcorner{k}}}:\mathbb{R}^{n(T+1)+n_{i}T}\times\mathbb{R}^{d}\rightarrow\mathbb{R}^{n_{\urcorner{k}}},\forall\beta\in[n_{\urcorner{k}}]$, as shown below:
$C_{k}$ $:=\{(\textbf{x},\textbf{u}):g_{\beta,k}(\textbf{x},\textbf{u})\leq 0,\forall\beta\in[n_{k}]\},$
$C_{\urcorner{k}}(\theta^{\star})$ $:=\{(\textbf{x},\textbf{u}):g_{\beta,{\urcorner{k}}}(\textbf{x},\textbf{u},\theta^{\star})\leq 0,\forall\beta\in[n_{\urcorner{k}}]\},$
For each $t\in[0,T-1]$, let $\pi_{t}:\mathbb{R}^{n_{o}(t+1)}\rightarrow\mathbb{R}^{n_{i}}$ be a causal output feedback controller which maps realized outputs $y_{0,t}$ to controls $u_{t}$, and let $\pi(\textbf{y}):=(\pi_{0}(y_{0}),\cdots,\pi_{T-1}(y_{0:T-1}))$. We define $\Pi:=\Pi(W,E,C_{k},C_{\urcorner{k}},\theta^{\star})$ to be the set of all causal output feedback laws $\pi(\cdot)$ which generate trajectories that robustly satisfy the constraint sets $C_{k}$ and $C_{\urcorner{k}}(\theta^{\star})$ despite worst-case noise realizations in $W$ and $E$. We then define a cost map $J:\mathbb{R}^{(n+1)T}\times\mathbb{R}^{mT}\rightarrow\mathbb{R}$ and formulate the following robust control problem, which is generally computationally intractable due to the optimization over $\pi(\cdot)$:
$\min_{\pi(\cdot)}.\max_{\textbf{w},\textbf{e},\textbf{x},\textbf{u},\textbf{y}}\hskip 8.53581pt$ $J(\textbf{x},\textbf{u})$ (3a)
[…]
$\textbf{w}\in W^{T+1},\textbf{e}\in E^{T+1},$ (3c)
$\pi(\cdot)\in\Pi(W,E,C_{k},C_{\urcorner{k}},\theta^{\star}),$ (3d)
II-B System Level Synthesis (SLS) for Output Feedback
A common method for (3) is to find (i) nominal state and control trajectories $(\textbf{z},\textbf{v})\in\mathbb{R}^{n(T+1)+n_{i}T}$ that minimize cost in the noise-free case and (ii) linear output feedback controllers to reject noise [15, 12, 16]. Equivalently, we assume each output feedback law $\pi_{t}(\cdot)$ for $t\in[0,T-1]$ has the form:
$u_{t}=\pi_{t}(y_{0:t})$ $=v_{t}+\sum_{\tau=0}^{t}K_{t,\tau}(y_{\tau}-C_{\tau}z_{\tau}),$ (4)
where $K_{t,\tau}\in\mathbb{R}^{n_{i}\times n_{o}}$, $\forall t\in[0,T-1]$, $\tau\in[0,t]$. To write (4) compactly, let $\mathcal{A}:=\text{diag}\{A_{0},\cdots,A_{T-1}\}\in\mathbb{R}^{nT\times nT}$, $\mathcal{B}:=\text{diag}\{B_{0},\cdots,B_{T-1}\}\in\mathbb{R}^{nT\times n_{i}T}$, $\mathcal{C}:=\text{diag}\{C_{0},\cdots,C_{T-1}\}\in\mathbb{R}^{n_{o}T\times nT}$, and define $\mathcal{K}\in\mathbb{R}^{n_{i}T\times n_{o}T}$ as follows:
$\mathcal{K}$ $:=\begin{bmatrix}K_{0,0}&O&\cdots&O\\[-7.0pt] K_{1,0}&K_{1,1}&\ddots&O\\[-3.0pt] \vdots&\vdots&\ddots&\vdots\\[-7.0pt] K_{T-1,0}&K_{T-1,1}&\ddots&K_{T-1,T-1}.\end{bmatrix}$ (5)
Then (4) can be written more compactly as:
$\textbf{u}=\pi(y_{0:T})$ $=\textbf{v}+\mathcal{K}(\textbf{y}-\mathcal{C}\textbf{z}).$ (6)
The policy optimization in (3) can now be recast as an optimization over the nominal state-control trajectory $(\textbf{z},\textbf{v})$ and the output error feedback gain $\mathcal{K}$ in (5). These two optimizations may be decoupled, as in methods based on control contraction metrics (CCM) [16], or coupled, as in system-level synthesis (SLS)-based robust control designs [15, 12].
Output feedback SLS [15, 11] parameterizes the feedback $\mathcal{K}$ via a system response $\Phi$, enabling joint optimization over $\mathcal{K}$ and $(\textbf{z},\textbf{v})$. Concretely, define the state, control, and output error respectively by $\Delta\textbf{x}:=\textbf{x}-\textbf{z}$, $\Delta\textbf{u}:=\textbf{u}-\textbf{v}$, and $\Delta\textbf{y}:=\textbf{y}-\mathcal{C}\textbf{z}$. Since $(\textbf{z},\textbf{v})$ satisfy the noise-free version of (1), i.e., $z_{t+1}=A_{t}z_{t}+B_{t}v_{t}$ for all $t\in[T]$, the error dynamics are:
$\Delta x_{t+1}$ $=A_{t}\Delta x_{t}+B_{t}\Delta u_{t}+w_{t},\hskip 14.22636pt$ $\forall t\in[0,T-1],$ (7a)
$\Delta u_{t}$ $=\sum_{\tau=0}^{t}K_{t,\tau}(C_{\tau}\Delta x_{\tau}+e_{\tau}),$ $\forall t\in[0,T-1],$ (7b)
with $\Delta x_{0}=0$. We concatenate (7) across times $t\in[T]$ by introducing the downshift operator $\mathcal{Z}$, given by:
$\mathcal{Z}$ $:=\begin{bmatrix}O_{n\times n(T-1)}&O_{n\times n}\\ I_{n(T-1)}&O_{n(T-1)\times n}\end{bmatrix}\in\mathbb{R}^{nT\times nT}.$ (8)
This enables us to rewrite (7) as:
$\Delta\textbf{x}$ $=\mathcal{Z}\mathcal{A}\Delta\textbf{x}+\mathcal{Z}\mathcal{B}\Delta\textbf{u}+\textbf{w},$ (9a)
$\Delta\textbf{u}$ $=\mathcal{K}(\mathcal{C}\Delta\textbf{x}+\textbf{e}).$ (9b)
Rearranging terms, we can rewrite (9) as:
$\begin{bmatrix}\Delta\textbf{x}\\ \Delta\textbf{u}\end{bmatrix}$ $=\begin{bmatrix}\Phi_{xw}&\Phi_{xe}\\ \Phi_{uw}&\Phi_{ue}\end{bmatrix}\begin{bmatrix}\textbf{w}\\ \textbf{e}\end{bmatrix}:=\Phi\begin{bmatrix}\textbf{w}\\ \textbf{e}\end{bmatrix}$, (10)
where the system response $\Phi\in\mathbb{R}^{(n(T+1)+n_{i}T)\times(n+n_{o})T}$ is block-wise defined from $\mathcal{Z},\mathcal{A},\mathcal{B},\mathcal{C}$, and $\mathcal{K}$, as follows:
$\Phi_{xw}$ $:=(I-\mathcal{Z}\mathcal{A}-\mathcal{Z}\mathcal{B}\mathcal{K}\mathcal{C})^{-1}$ (11a)
$\Phi_{xe}$ $:=(I-\mathcal{Z}\mathcal{A}-\mathcal{Z}\mathcal{B}\mathcal{K}\mathcal{C})^{-1}\mathcal{Z}\mathcal{B}\mathcal{K}$ (11b)
$\Phi_{uw}$ $:=\mathcal{K}\mathcal{C}(I-\mathcal{Z}\mathcal{A}-\mathcal{Z}\mathcal{B}\mathcal{K}\mathcal{C})^{-1}$ (11c)
$\Phi_{ue}$ $:=\mathcal{K}\mathcal{C}(I-\mathcal{Z}\mathcal{A}-\mathcal{Z}\mathcal{B}\mathcal{K}\mathcal{C})^{-1}\mathcal{Z}\mathcal{B}\mathcal{K}+\mathcal{K}.$ (11d)
Conversely, we also observe that:
$\mathcal{K}$ $=\Phi_{ue}-\Phi_{uw}\Phi_{xw}^{-1}\Phi_{xe}.$ (12)
Note that [15] proves that the closed-loop system response $\Phi$ characterizes all causal output feedback laws $\mathcal{K}$:
Proposition 1
[15, Prop. 1] If there exists a lower triangular block matrix $\mathcal{K}$ satisfying (9), then $\Phi_{xw}$, $\Phi_{xe}$, $\Phi_{uw}$, and $\Phi_{ue}$, as computed by (11) satisfy the affine equalities:
$\begin{bmatrix}I-Z\mathcal{A}&-Z\mathcal{B}\end{bmatrix}\Phi$ $=\begin{bmatrix}I&O\end{bmatrix}$, (13a)
$\Phi\begin{bmatrix}I-Z\mathcal{A}\\ -\mathcal{C}\end{bmatrix}$ $=\begin{bmatrix}I\\ O\end{bmatrix}$, (13b)
$\Phi_{xw},\Phi_{xe},\Phi_{uw},\Phi_{ue}$ block lower-triangular as in (5) (13c)
Conversely, if $\Phi_{xw}$, $\Phi_{xe}$, $\Phi_{uw}$, and $\Phi_{ue}$ satisfy (13), then $\mathcal{K}$, as computed by (12), is block lower-triangular.
Remark 1
If the dynamics (1a) are nonlinear, then (13), with each $A_{t}$, $B_{t}$ replaced by the Jacobian linearization of $f_{t}$ about given linearization points, would parameterize a subset of all causal affine output error feedback laws.
[…]
$\tilde{g}_{\beta,{\urcorner{k}}}(\textbf{z},\textbf{v},\Phi,\theta^{\star})$ $:=\max_{\begin{subarray}{c}\textbf{w}\in W^{T+1}\\ \textbf{e}\in E^{T+1}\end{subarray}}g_{\beta,{\urcorner{k}}}\left(\begin{bmatrix}\textbf{z}\\ \textbf{v}\end{bmatrix}+\Phi\begin{bmatrix}\textbf{w}\\ \textbf{e}\end{bmatrix},\theta^{\star}\right).$
We then stack constraints $\tilde{g}_{\beta,k}$ (resp., $\tilde{g}_{\beta,{\urcorner{k}}}$) across indices $\beta$ to form the vector-valued maps $\tilde{\textbf{g}}_{k}$ (resp., $\tilde{\textbf{g}}_{\urcorner{k}}$). We also define $\tilde{\textbf{h}}$ to capture the dynamics $z_{t+1}=A_{t}z_{t}+B_{t}v_{t}\ \forall t\in[T]$, as well as (13), as $\tilde{\textbf{h}}(\cdot)=0$. We then pose the following optimization program, which presents a feasible solution to (3) [12, Prop. 4]:
$\min_{\textbf{z},\textbf{v},\Phi}\hskip 14.22636pt$ $\tilde{J}(\textbf{z},\textbf{v},\Phi)$ (14a)
$\tilde{\textbf{h}}(\textbf{z},\textbf{v},\Phi)=\textbf{0},$ (14b)
$\tilde{\textbf{g}}_{k}(\textbf{z},\textbf{v},\Phi)\leq\textbf{0},$ (14c)
$\tilde{\textbf{g}}_{\urcorner{k}}(\textbf{z},\textbf{v},\Phi,\theta^{\star})\leq\textbf{0}.$ (14d)
Remark 2
If the dynamics (1a) are nonlinear, then $\tilde{\textbf{h}}$ must be defined to encode $z_{t+1}=f_{t}(z_{t},u_{t})$ for all $t\in[0,T]$, as well as (13), with each $A_{t},B_{t}$ replaced by the Jacobian linearization of $f_{t}$. Moreover, $\tilde{\textbf{g}}_{k}$ and $\tilde{\textbf{g}}_{\urcorner{k}}$ must be defined while accounting for the linearization error related to each $f_{t}$.
Remark 3
For control schemes where an output feedback law $\bar{\mathcal{K}}$ is designed independently of the optimal nominal trajectory (e.g., CCM methods), we can substitute $\mathcal{K}=\bar{\mathcal{K}}$ into (11), and append the resulting equalities (defining the resulting system response $\Phi$) to the constraint set of (14).
Given a parameter value $\theta\in\mathbb{R}^{d}$, let the set of corresponding safe (resp., unsafe) nominal trajectories and system responses, given below by $\mathcal{S}(\theta)$ (resp., $\mathcal{A}(\theta)$), be defined by:
$\mathcal{S}(\theta)$ $:=\{(\textbf{z},\textbf{v},\Phi):\tilde{\textbf{g}}_{\urcorner{k}}(\textbf{z},\textbf{v},\Phi,\theta)\leq\textbf{0}\},$ (15)
$\mathcal{A}(\theta)$ $:=\mathcal{S}(\theta)^{c}.$ (16)
In particular, $\mathcal{S}(\theta)$ contains the feasible set of (14).
Remark 4
$\tilde{g}_{\beta,k}(\cdot)$ and $\tilde{g}_{\beta,{\urcorner{k}}}(\cdot)$ can be written in closed form for certain constraint functions $g_{\beta,k}(\cdot)$ and $g_{\beta,{\urcorner{k}}}(\cdot)$ and disturbance sets $W$ and $E$ [12, Prop. 4], [13, Prop III.3].
II-C Demonstration Generation Process
We describe how a set of $D$ input-output trajectory demonstrations, denoted $\mathcal{D}:=\{(\textbf{u}^{({d})},\textbf{y}^{({d})})\in\mathbb{R}^{n_{i}T}\times\mathbb{R}^{n_{o}T}:d\in[D]\}$ are generated for constraint learning. First, the demonstrator solves (14) to compute a locally-optimal nominal trajectory $(z^{\star},v^{\star})$ and system response $\Phi^{\star}$, respectively, and computes the corresponding output feedback law $\mathcal{K}^{\star}$ by substituting $\Phi=\Phi^{\star}$ into (12). Then, for each $d\in[D]$, the demonstrator generates the input-output trajectory $(\textbf{u}^{({d})},\textbf{y}^{({d})})\in\mathbb{R}^{n_{i}T}\times\mathbb{R}^{n_{o}T}$, by unrolling the dynamics (1a), and applying output map (1b) and output feedback law (4):
$x_{t+1}^{({d})}$ $:=A_{t}x_{t}^{({d})}+B_{t}u_{t}^{({d})}+w_{t}^{({d})}$ (17a)
$y_{t}^{({d})}$ $=C_{t}x_{t}^{({d})}+e_{t}^{({d})}$ (17b)
$u_{t}^{({d})}$ $=v_{t}^{\star}+\sum_{\tau=0}^{t}K_{t,\tau}^{\star}(y_{\tau}^{({d})}-Cz_{\tau}^{\star}).$ (17c)
Above, $(\textbf{w}^{({d})},\textbf{e}^{({d})})\in W^{T+1}\times E^{T+1}$ describes the dynamics and output noise realizations for the $d$-th input-output trajectory, for each $d\in[D]$. Finally, the constraint learner receives the perturbed demonstration set $\tilde{D}:=\{(\tilde{\textbf{u}}^{({d})},\tilde{\textbf{y}}^{({d})})\in\mathbb{R}^{n_{i}T}\times\mathbb{R}^{n_{o}T}:d\in[D]\}$, with:
$\tilde{\textbf{u}}^{({d})}=\textbf{u}^{({d})}+\bm{\delta}_{u}^{({d})},\quad\quad\tilde{\textbf{y}}^{({d})}=\textbf{y}^{({d})}+\bm{\delta}_{y}^{({d})},$ (18)
where $(\bm{\delta}_{u}^{({d})},\bm{\delta}_{y}^{({d})})\in\mathbb{R}^{n_{i}T}\times\mathbb{R}^{n_{o}T}$ represent possible transmission errors that may corrupt the constraint learner’s reception of the generated input-output demonstrations.
Remark 5
Above, we considered multiple input-output trajectories generated from different rollouts of the same nominal trajectory using the same output feedback law. However, our methods also extend to learn constraints from multiple nominal trajectories and feedback laws.
II-D Problem Statement
We aim to infer the unknown constraint parameter $\theta^{\star}$ from the perturbed input-output demonstration set $\tilde{D}$, as described above, as well as knowledge of the system equations (1), costs $J$ and $\tilde{J}$, and constraints $g_{\beta,k}$, $g_{\beta,{\urcorner{k}}}$, $\tilde{g}_{\beta,k}$, and $\tilde{g}_{\beta,{\urcorner{k}}}$. We also aim to learn nominal trajectories and output feedback laws that are guaranteed to lie within the safe set $\mathcal{S}(\theta)$ despite worst-case dynamics and output noise realizations.
III Methods
We present our method for learning the nominal trajectory, output feedback law, and unknown constraints (Sec. III-A) and prove that under zero transmission noise and mild conditions, the recovered values are accurate (Sec. III-B). We also perform sensitivity analysis to prove that learning error grows linearly with transmission noise magnitude (Sec. III-C).
III-A Recovering the Nominal Trajectory, Output Feedback, and Constraint Parameters
First, we formulate a linear least-squares (LLS) program to infer the demonstrator’s output feedback policy $\mathcal{K}^{\star}$ from the perturbed demonstration set $\tilde{D}$ received by the constraint learner. Concretely, from (17c) and (18), we obtain:
$\tilde{\textbf{u}}^{({d})}-\bm{\delta}_{u}^{({d})}=\textbf{v}^{\star}+\mathcal{K}^{\star}(\tilde{\textbf{y}}^{({d})}-\mathcal{C}\textbf{z}^{\star}-\bm{\delta}_{y}^{({d})}),$ (19)
$\Rightarrow\$$(\tilde{\textbf{u}}^{({d})}-\tilde{\textbf{u}}^{({d-1})})-(\bm{\delta}_{u}^{({d})}-\bm{\delta}_{u}^{({d-1})})$ (20)
$\hskip 14.22636pt=\mathcal{K}^{\star}\big((\tilde{\textbf{y}}^{({d})}-\tilde{\textbf{y}}^{({d-1})})-(\bm{\delta}_{y}^{({d})}-\bm{\delta}_{y}^{({d-1})})\big)$
We define $\tilde{\textbf{U}}\in\mathbb{R}^{n_{i}T\times(D-1)}$, $\tilde{\textbf{Y}}\in\mathbb{R}^{n_{o}T\times(D-1)}$ and compute $\tilde{\mathcal{K}}$ as:
$\tilde{\textbf{U}}$ $:=\begin{bmatrix}\tilde{\textbf{u}}^{({2})}-\tilde{\textbf{u}}^{({1})}&\cdots&\tilde{\textbf{u}}^{({D})}-\tilde{\textbf{u}}^{({D-1})}\end{bmatrix}$, (21)
$\tilde{\textbf{Y}}$ $:=\begin{bmatrix}\tilde{\textbf{y}}^{({2})}-\tilde{\textbf{y}}^{({1})}&\cdots&\tilde{\textbf{y}}^{({D})}-\tilde{\textbf{y}}^{({D-1})}\end{bmatrix}$, (22)
$\tilde{\mathcal{K}}:=\tilde{\textbf{U}}\tilde{\textbf{Y}}^{\dagger}$. (23)
We then use (11) to compute the system response $\tilde{\Phi}$ corresponding to $\tilde{\mathcal{K}}$.
Next, to infer the nominal trajectories $(\textbf{z}^{\star},\textbf{v}^{\star})$, we stack the equalities $z_{t+1}^{\star}=A_{t}z_{t}^{\star}+B_{t}v_{t}^{\star}$ across $t\in[T]$ and rearrange (19) to obtain that, for each $d\in[D]$:
$(I-\mathcal{Z}\mathcal{A})\textbf{z}^{\star}-\mathcal{Z}\mathcal{B}\textbf{v}^{\star}$ $=\textbf{0},$
$-\mathcal{K}^{\star}\mathcal{C}\textbf{z}^{\star}+\textbf{v}^{\star}$ $=(\tilde{\textbf{u}}^{({d})}-\mathcal{K}^{\star}\tilde{\textbf{y}}^{({d})})-(\bm{\delta}_{u}^{({d})}-\mathcal{K}^{\star}\bm{\delta}_{y}^{({d})}).$
We then estimate $(\textbf{z}^{\star},\textbf{v}^{\star})$ from $\tilde{\mathcal{K}}$ as follows [^2] For the case of nonlinear dynamics, see Remark 6.:
$\begin{bmatrix}\tilde{\textbf{z}}\\ \tilde{\textbf{v}}\end{bmatrix}$ $=\frac{1}{d}\sum_{d=1}^{D}\begin{bmatrix}I-\mathcal{Z}\mathcal{A}&-\mathcal{Z}\mathcal{B}\\ -\tilde{\mathcal{K}}\mathcal{C}&I\end{bmatrix}^{\dagger}\begin{bmatrix}\textbf{0}\\ \tilde{\textbf{u}}^{({d})}-\tilde{\mathcal{K}}\tilde{\textbf{y}}^{({d})}\end{bmatrix}$. (24)
Above, we average over the terms $\{\tilde{\textbf{u}}^{({d})}-\tilde{\mathcal{K}}\tilde{\textbf{y}}^{({d})}:d\in[D]\}$ to reduce the impact of the transmission errors $\{\bm{\delta}_{u}^{({d})},\bm{\delta}_{y}^{({d})}:d\in[D]\}$ on the recovery of $(\textbf{z}^{\star},\textbf{v}^{\star})$.
Finally, we infer the unknown constraint parameter $\theta^{\star}$. For notational ease, given $\textbf{z},\textbf{v},\Phi$, we define $\bm{\eta}:=(\textbf{z},\textbf{v},\Phi)$. We also define $\bm{\nu}_{k}$, $\bm{\lambda}_{k}$, and $\bm{\lambda}_{\urcorner{k}}$ to denote the Lagrange multipliers associated with (14b), (14c), and (14d), respectively. The KKT conditions of (14) are then given by:
$\textbf{h}(\bm{\eta})=0,\hskip 14.22636pt\textbf{g}_{k}(\bm{\eta})\leq 0,\hskip 14.22636pt\textbf{g}_{\urcorner{k}}(\bm{\eta},\theta)\leq 0,$ (25a)
$\bm{\lambda}_{k},\bm{\lambda}_{\urcorner{k}}\geq 0,$ (25b)
$\bm{\lambda}_{k}\odot\textbf{g}_{k}(\bm{\eta})=0,\hskip 14.22636pt\bm{\lambda}_{\urcorner{k}}\odot\textbf{g}_{\urcorner{k}}(\bm{\eta},\theta)=0,$ (25c)
$\nabla_{\bm{\eta}}J(\bm{\eta})+\bm{\lambda}_{k}^{\top}\nabla_{\bm{\eta}}\textbf{g}_{k}(\bm{\eta})$ (25d)
$\hskip 5.69054pt+\bm{\lambda}_{\urcorner{k}}^{\top}\nabla_{\bm{\eta}}\textbf{g}_{\urcorner{k}}(\bm{\eta},\theta)+\bm{\nu}^{\top}\nabla_{\bm{\eta}}\textbf{h}(\bm{\eta})=0.$
We denote by $\text{KKT}(\bm{\eta})$ the set of all $(\theta,\bm{\lambda}_{k},\bm{\lambda}_{\urcorner{k}},\bm{\nu})$ satisfying (25). Since the true nominal trajectory and output response computed by the demonstrator, which we denote by $\bm{\eta}^{\star}:=(\textbf{z}^{\star},\textbf{v}^{\star},\Phi^{\star})$, minimizes (14), there exist Lagrange multiplier values $\bm{\lambda}_{k}^{\star},\bm{\lambda}_{\urcorner{k}}^{\star},\bm{\nu}^{\star}$ such that $(\theta^{\star},\bm{\lambda}_{k}^{\star},\bm{\lambda}_{\urcorner{k}}^{\star},\bm{\nu}^{\star})$ solves the following feasibility problem with $\bm{\eta}=\bm{\eta}^{\star}$:
find $\theta,\bm{\lambda}_{k},\bm{\lambda}_{\urcorner{k}},\bm{\nu},$ (26a)
s.t. $(\theta,\bm{\lambda}_{k},\bm{\lambda}_{\urcorner{k}},\bm{\nu})\in\text{KKT}(\bm{\eta}).$ (26b)
Conversely, solutions to (26) with $\bm{\eta}=\bm{\eta}^{\star}$ describe parameter values $\theta$ consistent with the optimality of $\bm{\eta}^{\star}$ with respect to (14), and thus inform the true parameter value $\theta^{\star}$. For convenience, given any $\bm{\eta}=(\textbf{z},\textbf{v},\Phi)$, we denote by $\mathcal{F}(\bm{\eta})\subseteq\mathbb{R}^{d}$ the set of KKT-compatible parameter values, i.e.,
$\mathcal{F}(\bm{\eta})$ $:=\{\theta\in\mathbb{R}^{d}:\exists\ \bm{\lambda}_{k},\bm{\lambda}_{\urcorner{k}},\bm{\nu}$ (27)
 $\hskip 56.9055pt\text{ s.t. }(\theta,\bm{\lambda}_{k},\bm{\lambda}_{\urcorner{k}},\bm{\nu})\in\text{KKT}(\bm{\eta}).\}$
III-B Learning Guarantees Under Zero Transmission Error
We now establish sufficient conditions under which $\bm{\eta}^{\star}$, the ground truth nominal trajectories and system response, can be accurately recovered.
Assumption 1
$\tilde{\textbf{Y}}$ is full row rank and $\Gamma$ (28) is invertible:
$\Gamma:=\begin{bmatrix}I-\mathcal{Z}\mathcal{A}&-\mathcal{Z}\mathcal{B}\\ -\mathcal{K}^{\star}\mathcal{C}&I\end{bmatrix}\in\mathbb{R}^{(n+n_{i})T\times(n+n_{i})T}.$ (28)
In words, Assumption 1 states that the provided demonstrations $\tilde{\textbf{Y}}$ are sufficiently rich for recovering $\mathcal{K}^{\star}$, and that $\mathcal{K}^{\star}$ satisfies a mild regularity condition. In particular, if $\bm{\delta}_{u}^{({d})}$ and $\bm{\delta}_{y}^{({d})}$ are i.i.d. across indices $d\in[D]$, then $\tilde{\textbf{Y}}$ has full row rank with zero mean whenever $D\geq n_{o}T+1$. Thm. 1 below shows that in the absence of transmission noise, Assumption 1 is sufficient to guarantee the accurate recovery of $\bm{\eta}^{\star}$.
Theorem 1
[…]
Proof:
When $(\bm{\delta}_{u}^{({d})},\bm{\delta}_{y}^{({d})})=\textbf{0},\forall d\in[D]$, (20) becomes:
$\tilde{\textbf{u}}^{({d})}-\tilde{\textbf{u}}^{({d-1})}$ $=\mathcal{K}^{\star}(\tilde{\textbf{y}}^{({d})}-\tilde{\textbf{y}}^{({d-1})}),$ (29)
which can be stacked across all $d\in[2,D]$ as $\tilde{\textbf{U}}=\mathcal{K}^{\star}\tilde{\textbf{Y}}$. Since $\tilde{\textbf{Y}}$ has full row rank, $\tilde{\textbf{Y}}\tilde{\textbf{Y}}^{\dagger}=I_{n_{o}T}$, so $\mathcal{K}^{\star}=\tilde{\textbf{U}}\tilde{\textbf{Y}}^{\dagger}$. Together with (23) and (11), this confirms that $\mathcal{K}^{\star}=\tilde{\mathcal{K}}$ and $\Phi^{\star}=\tilde{\Phi}$. Then, from (24) and preceding equations, we have under Assumption 1 that $(\tilde{\textbf{z}},\tilde{\textbf{v}})=(\textbf{z}^{\star},\textbf{v}^{\star})$. ∎
Remark 6
If the dynamics (1a) are nonlinear, we can still recover the output feedback law via (23) (since the output map is assumed LTV). However, we would apply nonlinear regression techniques to fit the input-output demonstrations, while enforcing the dynamics $z_{t+1}^{\star}=f_{t}(z_{t}^{\star},v_{t}^{\star})$, instead of applying the procedure described in Sec. III-A. Although the nominal trajectories would no longer be guaranteed to be accurately recovered (in the sense of Thm. 1), we find that in simulation, we still recover the true nominal trajectories with minimal error when enough demonstrations are provided.
We next define the set of guaranteed safe (resp., unsafe) nominal trajectories and output responses, denoted $\mathcal{G}_{s}(\bm{\eta}^{\star})$ (resp., $\mathcal{G}_{\urcorner{s}}(\bm{\eta}^{\star})$), to consist of nominal trajectories and output responses which are safe (resp., unsafe) with respect to all $\theta\in\mathcal{F}(\bm{\eta}^{\star})$. The set of guaranteed safe nominal trajectories and output feedback laws can then be used to generate motion plans downstream that will remain safe despite even worst-case noise realizations. Concretely, define: [^3] We remark that $\mathcal{G}_{\urcorner{s}}(\bm{\eta}^{\star})\neq\mathcal{G}_{s}(\bm{\eta}^{\star})^{c}$, i.e., a given nominal trajectory and system response may be neither guaranteed safe nor guaranteed unsafe.
$\mathcal{G}_{s}(\bm{\eta}^{\star})$ $:=\bigcap_{\theta\in\mathcal{F}(\bm{\eta}^{\star})}\big\{\bm{\eta}^{\prime}:\textbf{g}(\bm{\eta}^{\prime},\theta)\leq 0\big\}=\bigcap_{\theta\in\mathcal{F}(\bm{\eta}^{\star})}\mathcal{S}(\theta),$
$\mathcal{G}_{\urcorner{s}}(\bm{\eta}^{\star})$ $:=\bigcap_{\theta\in\mathcal{F}(\bm{\eta}^{\star})}\big\{\bm{\eta}^{\prime}:\textbf{g}(\bm{\eta}^{\prime},\theta)\leq 0\big\}^{c}=\bigcap_{\theta\in\mathcal{F}(\bm{\eta}^{\star})}\mathcal{A}(\theta).$
Below, Thm. 2 states that $\mathcal{G}_{s}(\bm{\eta}^{\star})$ is guaranteed to inner-approximate (i.e., conservatively estimate) $\mathcal{S}(\theta^{\star})$, the set of nominal trajectories and system responses $\bm{\eta}$ that are safe with respect to the ground truth inequality constraints. Similarly, $\mathcal{G}_{\urcorner{s}}(\bm{\eta}^{\star})$ always inner-approximates $\mathcal{A}(\theta^{\star})$.
Theorem 2
$\mathcal{G}_{s}(\bm{\eta}^{\star})\subseteq\mathcal{S}(\theta^{\star})$ and $\mathcal{G}_{\urcorner{s}}(\bm{\eta}^{\star})\subseteq\mathcal{A}(\theta^{\star})$.
Proof:
By the definitions of $\mathcal{G}_{s}(\bm{\eta}^{\star})$ and $\mathcal{G}_{\urcorner{s}}(\bm{\eta}^{\star})$, it suffices to prove $\theta^{\star}\in\mathcal{F}(\bm{\eta}^{\star})$. Since $\bm{\eta}^{\star}$ solves (14), there exist Lagrange multipliers $\bm{\lambda}_{k},\bm{\lambda}_{\urcorner{k}},\bm{\nu}$ such that $(\theta^{\star},\bm{\lambda}_{k},\bm{\lambda}_{\urcorner{k}},\bm{\nu})\in\text{KKT}(\bm{\eta}^{\star})$. Thus, by definition of $\mathcal{F}(\bm{\eta}^{\star})$, we have $\theta^{\star}\in\mathcal{F}(\bm{\eta}^{\star})$, so $\mathcal{G}_{s}(\bm{\eta}^{\star})\subseteq\mathcal{S}(\theta^{\star})$ and $\mathcal{G}_{\urcorner{s}}(\bm{\eta}^{\star})\subseteq\mathcal{A}(\theta^{\star})$. ∎
Remark 7
With respect to many real-world constraints (e.g., collision-avoidance), (26) can be reformulated as a mixed-integer linear program (MILP) or a mixed-integer bilinear program (MIBLP) [1, Sec. 4B, 4D], and solved via off-the-shelf solvers (e.g., Gurobi) [17].
III-C Sensitivity Analysis With Respect to Transmission Error
We now prove that when the transmission noise $\{(\bm{\delta}_{u}^{({d})},\bm{\delta}_{y}^{({d})}):d\in[D]\}$ is nonzero, the error in the recovered output feedback $\mathcal{K}$ and nominal trajectory $(\textbf{z},\textbf{v})$ scale at most linearly in the noise magnitudes $\|\bm{\delta}_{u}^{({d})}\|_{2}$ and $\|\bm{\delta}_{y}^{({d})}\|_{2}$.
Theorem 3
Suppose Assumption 1 holds, and $\max_{d\in[D]}\max\{\|\bm{\delta}_{u}^{({d})}\|_{2},\|\bm{\delta}_{y}^{({d})}\|_{2}\}<\epsilon$ for some $\epsilon>0$. Define $\textbf{Y}:=\begin{bmatrix}\textbf{y}^{({2})}-\textbf{y}^{({1})}&\cdots&\textbf{y}^{({D})}-\textbf{y}^{({D-1})}\end{bmatrix}$, $\rho_{1}=\sqrt{D-1}\|\textbf{Y}^{\dagger}\|_{2}$, $\rho_{2}=\sqrt{D-1}\|\Gamma\|_{2}\|\mathcal{C}\|_{2}$, $\rho_{3}=\sqrt{D-1}\|\Gamma\|_{2}\|\textbf{Y}^{\dagger}\|_{2}$, and $\rho_{4}=\|\Gamma\|_{2}\|\mathcal{K}^{\star}\|_{2}.$ Let $(\delta\textbf{z},\delta\textbf{v},\delta K):=(\tilde{\textbf{z}}-\textbf{z}^{\star},\tilde{\textbf{v}}-\textbf{v}^{\star},\tilde{\mathcal{K}}-\mathcal{K}^{\star})$. Then:
$\|\delta\mathcal{K}\|_{2}$ $\leq\rho_{1}\big(\|\tilde{\mathcal{K}}\|_{2}+1\big)\epsilon,$
$\left\|\begin{bmatrix}\delta\textbf{z}\\ \delta\textbf{v}\end{bmatrix}\right\|_{2}$ $\leq\left(\rho_{2}\left\|\begin{bmatrix}\tilde{\textbf{z}}\\ \tilde{\textbf{v}}\end{bmatrix}\right\|_{2}+\rho_{3}\|\tilde{\textbf{y}}^{({1})}\|_{2}(\|\tilde{\mathcal{K}}\|_{2}+1)+\rho_{4}\right)\epsilon.$
Proof:
[…]
IV Experiments
To evaluate our method, we perform constraint learning on simulated 4D double integrator, 4D unicycle, nonlinear 6D quadcopter, and linearized 12D quadcopter dynamics [19].
IV-A Experiment Setup
We encode polytopic collision-avoidance constraints using intersections of unions of half-spaces, as shown below:
$C_{\urcorner{k}}(\theta)$ $=\bigwedge_{\alpha\in[\overline{N}_{c}]}\bigvee_{\beta\in[N_{c}]}\big\{A_{\alpha,\beta}(\theta)(\textbf{x},\textbf{u})\leq b_{\alpha,\beta}(\theta)\big\},$
where $A_{\alpha,\beta}(\theta)$ and $b_{\alpha,\beta}$ are dimension-compatible $\forall\alpha,\beta$.
To define agent costs $J$, given a state trajectory $\textbf{x}\in\mathbb{R}^{nT}$, let the vector $p_{t}$ and the scalars $p_{x,t},p_{y,t}$ denote components of x describing the position vector, $x$-coordinate and $y$-coordinate at time $t$, respectively. We then define:
$J^{({1})}(\textbf{x})$ $=\textstyle\sum_{t=0}^{T-1}\big(\|p_{t+1}-p_{t}\|_{2}^{2}-p_{x,t}\big),$ (31a)
$J^{({2})}(\textbf{x})$ $=\textstyle\sum_{t=0}^{T-1}\big(\|p_{t+1}-p_{t}\|_{2}^{2}-p_{x,t}-p_{y,t}\big),$ (31b)
$J^{({3})}(\textbf{x})$ $=\textstyle\sum_{t=0}^{T-1}\Big(\|p_{t+1}-p_{t}\|_{2}^{2}+\frac{1}{T}\|p_{t}-p_{T}\|_{2}^{2}\Big).$ (31c)
Demonstrations in Figs. 2c and 2d-f were generated with $J^{({3})}$ and $J^{({2})}$, respectively; all others were generated using $J^{({1})}$.
In the Figs. 2c, 3a-b, 4a-c simulations, the set $W$ bounding the dynamics noise terms $w_{t}$ was defined to be $\infty$-norm balls of radii 0.05, 0.05, 0.05, 0.2, 0.07, and 0.05, respectively; in all other simulations, $W$ was set as the unit $\infty$-norm ball. In the Fig. 2 simulations, the set $V$ bounding the dynamics noise terms $v_{t}$ was set as the unit norm ball of radius 0.02; all other simulations used zero output noise.
As shown in Figs. 2-4, for each experiment, we fix a ground truth constraint set (shaded gray, with yellow boundaries), and use output SLS, control contraction metrics (CCM), or proportional–derivative (PD) control to generate nominal trajectories (green, with circles/stars for origins/goals), output feedback laws, and demonstration rollouts (blue) with corresponding reachable tubes and tube edges (magenta). We then apply the methods in Sec. III to learn the nominal trajectories, output feedback laws, and constraints (boundaries in black) from the given demonstrations.
For each experiment, we generate demonstrations either via (full) state feedback (i.e., $y_{t}=x_{t}$), or output feedback, wherein only the position vector is observed (i.e., $y_{t}=p_{t}$). For the baseline experiment (Fig. 2b), we use $D=20$ demonstrations; for all others, we set $D=100$. The simulations in Figs. 2-4 are generated with zero transmission noise; Fig. 1 shows that the learning accuracy of our method degrades gracefully as transmission error magnitude increases.
Figure 1: Error magnitude in the recovered nominal trajectories, nominal controls, output feedback, and constraint parameter as a function of the transmission error level. Demonstrations were generated using noise-corrupted double integrator dynamics and (full) state feedback SLS controllers. On average, our method enjoys high accuracy in learning a priori unknown parameters under low levels of transmission noise.
IV-B Simulation Results
Figure 2: Constraint learning from demonstrations generated via SLS with double integrator (a, c) and linear 12D quadcopter (d, e, f) dynamics. (e, f) provide views of (d) from different angles. (a, c-f) Our method accurately learns the true collision avoidance constraints across all simulations. (b) In contrast, the baseline [1] failed to accurately recover the true constraint; note the mismatch between the learned (black) and true (yellow) constraints.
SLS Controllers
We present simulation results evaluating our constraint recovery method (Figs. 2a, c-f) and comparing it to the baseline in [1] (Fig. 2b), which was designed to learn constraints from noise-free, fully observed state trajectories. We learn constraints from demonstrations generated by unrolling noise-corrupted double integrator (Fig. 2 a-c) or linearized 12D quadcopter (Fig. 2d-f) dynamics, and applying a state feedback (Fig. 2 a-b, d-f) or output feedback (Fig. 2c) SLS controller (Fig. 2c). The nominal trajectories and feedback laws underlying the input-output demonstration were jointly generated via (14). Across experiments in Figs. 2a, c-f, our methods, as described in (23)-(26), accurately learned the true nominal trajectories, output feedback laws, and constraint parameters underlying the input-output demonstrations. In contrast, the baseline [1] fails to learn the true underlying constraints (Fig. 2b).
Figure 3: Constraint learning from demonstrations produced via CCM with unicycle (a) and double integrator (b) dynamics. Our method accurately learns the true constraints for all simulations.
Control Contraction Metrics (CCM)-Based State Feedback Controllers
We learn constraints from demonstrations generated by unrolling noise-corrupted unicycle (Fig. 3a) or double integrator (Fig. 3b) dynamics with full state observation, using the CCM-based state feedback law in [16]. Concretely, we computed the error feedback law $\mathcal{K}$ prescribed in [16, Sec. III-B], and the associated system response $\Phi$ via (11). We then use (14) to compute the nominal trajectory for generating demonstrations, while fixing $\Phi$ fixed at the value derived from via CCM (see Remark 2). Across experiments in Fig. 3, by applying (23)-(26), we accurately learned the true nominal trajectories, output feedback laws, and constraint parameters underlying the demonstrations.
Figure 4: Constraint learning from demonstrations produced via CCM with (a) unicycle, (b) double integrator, and (c) nonlinear 6D quadcopter dynamics. Our method accurately learns the true constraints across all simulations.
Proportional-Derivative (PD) Controllers
We recover constraints from demonstrations generated using noise-corrupted unicycle (Fig. 4a), double integrator (Fig. 4b), or nonlinear 6D quadcopter (Fig. 4c) dynamics, by applying PD state feedback controllers. Similar to the CCM-based methods, we first compute the feedback policy $\mathcal{K}$ and the corresponding system response $\Phi$ via PD control techniques and (11). To obtain a nominal trajectory from which demonstrations are then generated, we solve (14) while holding $\Phi$ fixed at its value computed via the PD control method and (11). In the Fig. 4 experiments, our methods, as described by (23)-(26) accurately recovered the true nominal trajectories, output feedback laws, and constraint parameters that characterize the input-output demonstrations.
V Conclusion
We presented a novel IOC-based method for recovering unknown parametric constraints from locally optimal input-output demonstrations, generated by unrolling stochastic dynamics while applying an output feedback law designed using robust optimal control. We present theory and simulations to demonstrate that our method can accurately recover the demonstrator’s output feedback laws and constraints.